OpenStudy (anonymous):

s+2 and 1/6=4 and 2/3. what does s equal???

OpenStudy (anonymous):

Did you forget some variables?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

1/6 = 4 makes no sense...

OpenStudy (anonymous):

did you mean 1/6s = 4?

OpenStudy (anonymous):

and we figured this problem out ourselves thank you very much. :P

OpenStudy (anonymous):

I shall give you a medal then :)

OpenStudy (anonymous):

thank ya!!! :P lol

jhonyy9 (jhonyy9):

s+2 1/6 =4 2/3 s+(6*2+1)/6 =(3*4+2)/3 s+13/6 =14/3 (2s+13)/6 =(2*14)/6 2s+13=2*14 2s+13=28 2s=28-13 2s=15 s=15/2 s=7 1/2